OpenStudy (anonymous):

What is the inverse of f(x) = 2x − 8?

OpenStudy (anonymous):

switch the x and y's and solve for Y

OpenStudy (amistre64):

x = 2y-8...solve for y x+8 ---- = y should be the inverse 2

OpenStudy (anonymous):

f^1x=(x+8)/2